let A (2,5) and B(4, -1) be vertices of a triangle ABC. if the centroid of the triangle moves on the line 2x + 3y = 1, then locus of the vertex C is the line

Grade:11

1 Answers

praneethvikram
5 years ago
 given centroid lies on 2x + 3y =1
 
a ( 2,5 )  b (4 , -1)  let c be (m,n)
 
=> centroid will be  (6+m)/3 , (4+n)/3
 
it lies on 2x + 3y =1
substitute it
4 + 2m/3 + 4 + n =1
=> 2m/3 + n =-7
=> locus is  2m + 3n = -21